OpenStudy (anonymous):

write an equation for the circle with center(-1,6), containing (2,1)

OpenStudy (mertsj):

(x+1)^2+(y-6)^2=34

OpenStudy (anonymous):

how did you get that?? thanks

OpenStudy (anonymous):

only need the square of the radias, \((-1-2)^2+(6-1)^2=3^2+5^2=35\)

OpenStudy (anonymous):

oops i meant 34

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

square of the distance from (-1,6) to (2,1) since equation is \((x-h)^2+(y-k)^2=r^2\)